OpenStudy (anonymous):

write an explicit formula for the sequence 4, 1.5, -1, -3.5

OpenStudy (amistre64):

what have you tried?

OpenStudy (anonymous):

well i figured out how to write the recursive formula but I cant figure out the explicit

OpenStudy (amistre64):

show me the recurrsive

OpenStudy (anonymous):

\[A _{n}=A _{n-1}-2.5\]

OpenStudy (amistre64):

your missing a part, a recurrsive rule as the equation part, and also an A1 part. You need to be able to define where the sequence starts at.

OpenStudy (anonymous):

where A1 = 4

OpenStudy (amistre64):

good A1 = 4 A2 = A1 - 2.5 A3 = A2 - 2.5 A4 = A3 - 2.5 you agree that this is how it would progress recurrsively then right?

OpenStudy (amistre64):

if so, then lets start defining the An terms in just A1 A1 = 4 A2 = A1 - 2.5 = 4 - 2.5 A3 = A2 - 2.5 = 4 - 2.5 - 2.5 = 4 -2.5(2) A4 = A3 - 2.5 = 4 - 2.5 - 2.5 - 2.5 = 4 -2.5(3) ..................... or to clean it up some lets write them like this A1 = 4 - 2.5(0) A2 = 4 - 2.5(1) A3 = 4 -2.5(2) A4 = 4 -2.5(3) notice that the nth A follows a pattern: An = 4 - 2.5(n-1) seems to fit

OpenStudy (anonymous):

okay im kindof confused